# How to Adapt Control Barrier Functions? A Learning-Based Approach with Applications to a VTOL Quadplane

### v4

@@ I INTRODUCTION
Autonomous systems must operate while respecting safety constraints such as remaining within operational bounds or avoiding collisions. Control Barrier Functions have emerged as an effective tool for enforcing such constraints by ensuring forward invariance of a predefined safe set [1]. Enforcing CBF constraints for safety defines a class of optimization problems, typically Quadratic Programs and Model Predictive Control (MPC) formulations, that can be solved efficiently in real time [2, 3]. This property has made CBFs popular in a range of applications, from robotic navigation [4] to multi-agent coordination [5].
However, the definition of CBF constraints entails a class of functions called class $\mathcal{K}_{\infty}$ functions, which we refer to as CBF parameters. Finding or tuning the CBF parameters is particularly challenging for highly nonlinear and complex systems and becomes even more difficult when input constraints are present [6]. In practice, one can resort to fixing the CBF parameters a priori based on trial and error. However, using manually-tuned parameter values can lead to overly conservative motions or unsafe behavior when the environment or operating conditions change.
@@
Several recent works have recognized the need to adjust CBF parameters online. For instance, in [7, 8], these parameters are optimized to maintain controller feasibility at each time step; however, these approaches fall short of ensuring that safety is maintained for all time. Other methodologies leverage neural networks or reinforcement learning policies to dynamically adapt CBF parameters. Although these techniques have demonstrated promising performance in complex environments, their safety guarantees often rely on unrealistic assumptions, such as unbounded control authority [9] or perfect model predictions [10, 11, 12]. When those assumptions are relaxed, safety can no longer be guaranteed.
A critical factor in the reliability of learning-based adaptation and control is the treatment of uncertainty in neural network predictions. Neural network uncertainty can be broadly classified into two categories: aleatoric and epistemic [13]. Aleatoric uncertainty stems from inherent data stochasticity such as observation and process noise, whereas epistemic uncertainty arises from insufficient training data. Neglecting these uncertainties can lead to unreliable predictions, particularly when facing out-of-distribution inputs, which may lead to catastrophic failures in safety-critical applications.
Motivated by these findings, we propose a learning-based approach grounded in the previously described theoretical framework for online adaptation of CBF parameters. In particular, we extend our previous work, which employed Probabilistic Ensemble Neural Networks (PENNs) [14], and integrate it into our framework as one concrete solution for adapting CBF parameters. Our approach explicitly accounts for both epistemic and aleatoric uncertainties inherent in neural network predictions, ensuring that the adapted parameters consistently satisfy the theoretical conditions required for local safety guarantees.
@@ II-A Tangent Cone
Definition 1 (Tangent Cone [15]).
@@ II-B Safety Analysis of CBFs via Tangent Cone
Theorem 1 (Nagumo’s Theorem [16]).
@@ II-C Input Constrained CBFs
We extend the notion of CBFs to incorporate input constraints, yielding Input Constrained CBFs (ICCBFs) [6], a generalization of High Order CBFs (HOCBFs) [17].
@@
Definition 5 (Inner Safe Set [14]).
@@
Definition 6 (ICCBF [6]).
The function $h$ is a ICCBF [^3] In [6], $b_{r-1}$ is referred to as ICCBF; in this paper, for notational consistency with HOCBF [17], we designate $h$ as ICCBF, with all corresponding theoretical properties preserved. on $\mathcal{C}^{*}(\bm{\alpha})$ for System (2) if there exist class $\mathcal{K}$ functions $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{r}\}$ such that
@@ III METHOD OVERVIEW
Definition 7 (Locally Validated CBF Parameter [^5] It was also termed a locally valid parameter in [14].).
@@ IV PRIOR LEARNING-BASED APPROACHES TO CBF PARAMETER ADAPTATION
While the works in [9, 10, 11, 12] do not explicitly implement our locally validated CBF parameter adaptation framework, they fall into this category and have demonstrated promising results in reducing conservatism through online adaptation of CBF parameters using neural networks. However, these approaches do not sufficiently account for the inherent uncertainty in neural network predictions. Consequently, the safety of the system cannot be guaranteed, as the predicted CBF parameter $\tilde{\bm{\alpha}}^{*}$ out of (23) might not lie within the set of locally validated CBF parameters $\mathcal{A}$ given by Problem 2.
@@ V-A Probabilistic Ensemble Neural Network Model
We employ a Probabilistic Ensemble Neural Network (PENN) as the prediction model, denoted as $\bf F$}, to capture both prediction uncertainty and inherent output variability [18]. Building on the input $\bar{X}$ from Sec. IV, we augment it with the CBF parameter of interest $\tilde{\bm{\alpha}}$. Given this augmented input $X=[\bar{X}^{\top},\tilde{\bm{\alpha}}$$^{\top}$, the model outputs two quantities $Y=[\phi,\delta]^{\top}$: a safety margin $\phi\in\mathbb{R}$ and a performance metric $\delta\in\mathbb{R}$ (see Fig. 2b). Concretely, we define $\phi$ as the minimum of the ICCBF constraint function over the interval $[t,t+T]$:
@@
To capture epistemic uncertainty, we independently initialize each ensemble member with random weights $\bm{\theta}_{b}$. The resulting disagreement among ensemble predictions serves as a quantitative measure of epistemic uncertainty. Another notable advantage is that the forward pass of the PENN for prediction can be fully parallelized across ensemble members, which is beneficial for real-time implementation [19].
@@ V-B1 Verification Under Epistemic Uncertainty
In this step, we first evaluate the epistemic uncertainty associated with the PENN predictions. Specifically, we quantify the disagreement among ensemble members using the Jensen-Rényi divergence (JRD) with quadratic Rényi entropy, which admits a closed-form expression for Gaussian mixtures [20]. In essence, the JRD $D(X;{\bf E}_{1:B})\in\mathbb{R}^{\dim(Y)}$ measures how “spread out” or inconsistent the ensemble predictions are.
If the computed JRD exceeds a predefined threshold $D_{\textup{thr}}\in\mathbb{R}^{\dim(Y)}$, the corresponding input is deemed to be out-of-distribution and its prediction is discarded. This serves as our first “gate” to ensure that only high-confidence predictions interpolating within the training data, rather than extrapolating far beyond them, are used to determine the local validity of the CBF parameter.
@@ V-B2 Verification Under Aleatoric Uncertainty
We utilize the Conditional Value at Risk (CVaR) to assess the left-tail risk associated with $\hat{\phi}$ [21]. For a single ensemble member’s prediction $\hat{\phi}\sim{\bf E}_{b}(X)$, the left Value at Risk (VaR) at confidence level $\varepsilon$ is defined as
@@
Building on VaR, we define CVaR using an optimization formulation as proposed in [22]:
@@
To robustly account for aleatoric uncertainty, we treat the ensemble set $\mathfrak{E}=\{{\bf E}_{1},\ldots,{\bf E}_{B}\}$ as an ambiguity set and impose the following distributionally robust CVaR constraint [23]:
$\sup_{{\bf E}_{b}\in\mathfrak{E}}\textnormal{CVaR}{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\geq 0.$ (28)
This condition is equivalent to requiring that
@@ VI-A1 Aerodynamics Forces
To capture realistic airfoil stall effects at high angles of attack, the lift coefficient $C_{L}(\alpha)$ is chosen to be a smooth nonlinear function of $\alpha$ that blends a linear lift curve at small $|\alpha|$ with a flat-plate model at large $|\alpha|$ [24]. Concretely,
@@
Following [25], the aerodynamic forces in the body frame are given by
@@ VI-A2 Control Surface and Rotor Forces
Similarly, the total pitching moment, $M_{\textup{total}}$, is modeled following [25].
@@ VI-B Experimental Setup & Results
The test environment, illustrated in Fig. 4, features a scenario where the quadplane initiates in cruise mode with a high initial x-axis velocity, transitions to hover mode at an intermediate waypoint $(70,10)$ m, and lands at $(70,0)$ m while avoiding obstacles. The candidate ICCBF for collision avoidance is defined by the distance-based constraint: $\tilde{h}_{j}({\bm{x}})=(x-x_{\textup{obs},j})^{2}+(z-z_{\textup{obs},j})^{2}-(l_{\textup{robot}}+l_{\textup{obs},j})^{2}$, where $(x_{\textup{obs},j},z_{\textup{obs},j})$ and $l_{\textup{obs},j}$ denote the location and radius of the $j^{\textup{th}}$ obstacle, and $l_{\textup{robot}}$ is the quadplane’s safety radius. The safety margin $\phi\in Y$ is computed as the minimum across all obstacles. We use $r=2$ to construct the candidate ICCBF, consistent with the fact that $\tilde{h}_{j}$ has a relative degree of 2 for all four control inputs. As demonstrated in [6], the ICCBF thus trivially reduces to the HOCBF.
For training, we generate $22,050$ trajectory samples offline by varying the initial state $\bm{x}$} and the linear CBF parameters $\tilde{\bm{\alpha}}=(\tilde{\alpha}_{1},\tilde{\alpha}_{2})$ [^8] In this work, we use linear class $\mathcal{K}_{\infty}$ functions for simplicity.. The forward propagation interval $T$, representing the maximum simulation duration, is set to $15$ s to allow sufficient time for the system to land when feasible. In cases of infeasibility or collision, the safety margin $\phi$ is assigned a large negative value to clearly distinguish those from marginally feasible cases (i.e., where $0\leq\phi\ll 1$). We do not include additional environmental information in the neural network input $X$, given that obstacles are static in this test case. The PENN model $\bf F$} comprises $B=3$ ensemble models, each has 5 MLP layers with ReLU activation function. During training, 3% i.i.d. Gaussian noise is added to the input $X$, excluding the CBF parameters, to simulate observation and dynamics model uncertainties. Further implementation details follow [14].

### v5

@@ I INTRODUCTION
Autonomous systems must operate while respecting safety constraints such as remaining within operational bounds or avoiding collisions. Control Barrier Functions have emerged as an effective tool for enforcing such constraints by ensuring forward invariance of a predefined safe set [ames_control_2019]. Enforcing CBF constraints for safety defines a class of optimization problems, typically Quadratic Programs and Model Predictive Control (MPC) formulations, that can be solved efficiently in real time [ames_control_2014, zeng_safetycritical_2021]. This property has made CBFs popular in a range of applications, from robotic navigation [kim_visibility-aware_2025] to multi-agent coordination [zhang_gcbf_2025].
However, the definition of CBF constraints entails a class of functions called class $\mathcal{K}_{\infty}$ functions, which we refer to as CBF parameters. Finding or tuning the CBF parameters is particularly challenging for highly nonlinear and complex systems and becomes even more difficult when input constraints are present [agrawal_safe_2021]. In practice, one can resort to fixing the CBF parameters a priori based on trial and error. However, using manually-tuned parameter values can lead to overly conservative motions or unsafe behavior when the environment or operating conditions change.
@@
Several recent works have recognized the need to adjust CBF parameters online. For instance, in [zeng_safetycritical_2021_decay, zeng_enhancing_2021], these parameters are optimized to maintain controller feasibility at each time step; however, these approaches fall short of ensuring that safety is maintained for all time. Other methodologies leverage neural networks or reinforcement learning policies to dynamically adapt CBF parameters. Although these techniques have demonstrated promising performance in complex environments, their safety guarantees often rely on unrealistic assumptions, such as unbounded control authority [ma_learning_2022] or perfect model predictions [xiao_barriernet_2023, gao_online_2023, mohammad_soft_2025]. When those assumptions are relaxed, safety can no longer be guaranteed.
A critical factor in the reliability of learning-based adaptation and control is the treatment of uncertainty in neural network predictions. Neural network uncertainty can be broadly classified into two categories: aleatoric and epistemic [abdar_review_2021]. Aleatoric uncertainty stems from inherent data stochasticity such as observation and process noise, whereas epistemic uncertainty arises from insufficient training data. Neglecting these uncertainties can lead to unreliable predictions, particularly when facing out-of-distribution inputs, which may lead to catastrophic failures in safety-critical applications.
Motivated by these findings, we propose a learning-based approach grounded in the previously described theoretical framework for online adaptation of CBF parameters. In particular, we extend our previous work, which employed Probabilistic Ensemble Neural Networks (PENNs) [kim_learning_2025], and integrate it into our framework as one concrete solution for adapting CBF parameters. Our approach explicitly accounts for both epistemic and aleatoric uncertainties inherent in neural network predictions, ensuring that the adapted parameters consistently satisfy the theoretical conditions required for local safety guarantees.
@@ II-A Tangent Cone
Definition 1 (Tangent Cone [blanchini_set-theoretic_2008]).
@@ II-B Safety Analysis of CBFs via Tangent Cone
Theorem 1 (Nagumo’s Theorem [nagumo_uber_1942]).
@@ II-C Input Constrained CBFs
We extend the notion of CBFs to incorporate input constraints, yielding Input Constrained CBFs (ICCBFs) [agrawal_safe_2021], a generalization of High Order CBFs (HOCBFs) [xiao_control_2019].
@@
Definition 5 (Inner Safe Set [kim_learning_2025]).
@@
Definition 6 (ICCBF [agrawal_safe_2021]).
The function $h$ is a ICCBF [^3] In [agrawal_safe_2021], $b_{r-1}$ is referred to as ICCBF; in this paper, for notational consistency with HOCBF [xiao_control_2019], we designate $h$ as ICCBF, with all corresponding theoretical properties preserved. on $\mathcal{C}^{*}(\bm{\alpha})$ for System (2) if there exist class $\mathcal{K}$ functions $\bm{\alpha}=\{\alpha_{1},\ldots,\alpha_{r}\}$ such that
@@ III METHOD OVERVIEW
Definition 7 (Locally Validated CBF Parameter [^5] It was also termed a locally valid parameter in [kim_learning_2025].).
@@ IV PRIOR LEARNING-BASED APPROACHES TO CBF PARAMETER ADAPTATION
While the works in [ma_learning_2022, xiao_barriernet_2023, gao_online_2023, mohammad_soft_2025] do not explicitly implement our locally validated CBF parameter adaptation framework, they fall into this category and have demonstrated promising results in reducing conservatism through online adaptation of CBF parameters using neural networks. However, these approaches do not sufficiently account for the inherent uncertainty in neural network predictions. Consequently, the safety of the system cannot be guaranteed, as the predicted CBF parameter $\tilde{\bm{\alpha}}^{*}$ out of (23) might not lie within the set of locally validated CBF parameters $\mathcal{A}$ given by Problem 2.
@@ V-A Probabilistic Ensemble Neural Network Model
We employ a Probabilistic Ensemble Neural Network (PENN) as the prediction model, denoted as $\bf F$}, to capture both prediction uncertainty and inherent output variability [chua_deep_2018]. Building on the input $\bar{X}$ from Sec. IV, we augment it with the CBF parameter of interest $\tilde{\bm{\alpha}}$. Given this augmented input $X=[\bar{X}^{\top},\tilde{\bm{\alpha}}$$^{\top}$, the model outputs two quantities $Y=[\phi,\delta]^{\top}$: a safety margin $\phi\in\mathbb{R}$ and a performance metric $\delta\in\mathbb{R}$ (see Fig. 2b). Concretely, we define $\phi$ as the minimum of the ICCBF constraint function over the interval $[t,t+T]$:
@@
To capture epistemic uncertainty, we independently initialize each ensemble member with random weights $\bm{\theta}_{b}$. The resulting disagreement among ensemble predictions serves as a quantitative measure of epistemic uncertainty. Another notable advantage is that the forward pass of the PENN for prediction can be fully parallelized across ensemble members, which is beneficial for real-time implementation [kim_bridging_2023].
@@ V-B1 Verification Under Epistemic Uncertainty
In this step, we first evaluate the epistemic uncertainty associated with the PENN predictions. Specifically, we quantify the disagreement among ensemble members using the Jensen-Rényi divergence (JRD) with quadratic Rényi entropy, which admits a closed-form expression for Gaussian mixtures [wang_closed-form_2009]. In essence, the JRD $D(X;{\bf E}_{1:B})\in\mathbb{R}$ measures how “spread out” or inconsistent the ensemble predictions are.
If the computed JRD exceeds a predefined threshold $D_{\textup{thr}}\in\mathbb{R}$, the corresponding input is deemed to be out-of-distribution and its prediction is discarded. This serves as our first “gate” to ensure that only high-confidence predictions interpolating within the training data, rather than extrapolating far beyond them, are used to determine the local validity of the CBF parameter.
@@ V-B2 Verification Under Aleatoric Uncertainty
We utilize the Conditional Value at Risk (CVaR) to assess the left-tail risk associated with $\hat{\phi}$ [cai_evora_2024]. For a single ensemble member’s prediction $\hat{\phi}\sim{\bf E}_{b}(X)$, the left Value at Risk (VaR) at confidence level $\varepsilon$ is defined as
@@
Building on VaR, we define CVaR using an optimization formulation as proposed in [rockafellar_optimization_2000]:
@@
To robustly account for aleatoric uncertainty, we treat the ensemble set $\mathfrak{E}=\{{\bf E}_{1},\ldots,{\bf E}_{B}\}$ as an ambiguity set and impose the following distributionally robust CVaR constraint [rahimian_distributionally_2022]:
$\inf_{{\bf E}_{b}\in\mathfrak{E}}\textnormal{CVaR}_{\varepsilon}^{{\bf E}_{b}(X)}\left(\hat{\phi}\right)\geq 0.$ (28)
This condition guarantees that
@@ VI-A1 Aerodynamics Forces
To capture realistic airfoil stall effects at high angles of attack, the lift coefficient $C_{L}(\alpha)$ is chosen to be a smooth nonlinear function of $\alpha$ that blends a linear lift curve at small $|\alpha|$ with a flat-plate model at large $|\alpha|$ [willis_pitch_2022]. Concretely,
@@
Following [beard_small_2012], the aerodynamic forces in the body frame are given by
@@ VI-A2 Control Surface and Rotor Forces
Similarly, the total pitching moment, $M_{\textup{total}}$, is modeled following [beard_small_2012].
@@ VI-B Experimental Setup & Results
The test environment, illustrated in Fig. 4, features a scenario where the quadplane initiates in cruise mode with a high initial x-axis velocity, transitions to hover mode at an intermediate waypoint $(70,10)$ m, and lands at $(70,0)$ m while avoiding obstacles. The candidate ICCBF for collision avoidance is defined by the distance-based constraint: $\tilde{h}_{j}({\bm{x}})=(x-x_{\textup{obs},j})^{2}+(z-z_{\textup{obs},j})^{2}-(l_{\textup{robot}}+l_{\textup{obs},j})^{2}$, where $(x_{\textup{obs},j},z_{\textup{obs},j})$ and $l_{\textup{obs},j}$ denote the location and radius of the $j^{\textup{th}}$ obstacle, and $l_{\textup{robot}}$ is the quadplane’s safety radius. The safety margin $\phi\in Y$ is computed as the minimum across all obstacles. We use $r=2$ to construct the candidate ICCBF, consistent with the fact that $\tilde{h}_{j}$ has a relative degree of 2 for all four control inputs. As demonstrated in [agrawal_safe_2021], the ICCBF thus trivially reduces to the HOCBF.
For training, we generate $22,050$ trajectory samples offline by varying the initial state $\bm{x}$} and the linear CBF parameters $\tilde{\bm{\alpha}}=(\tilde{\alpha}_{1},\tilde{\alpha}_{2})$ [^8] In this work, we use linear class $\mathcal{K}_{\infty}$ functions for simplicity.. The forward propagation interval $T$, representing the maximum simulation duration, is set to $15$ s to allow sufficient time for the system to land when feasible. In cases of infeasibility or collision, the safety margin $\phi$ is assigned a large negative value to clearly distinguish those from marginally feasible cases (i.e., where $0\leq\phi\ll 1$). We do not include additional environmental information in the neural network input $X$, given that obstacles are static in this test case. The PENN model $\bf F$} comprises $B=3$ ensemble models, each has 5 MLP layers with ReLU activation function. During training, 3% i.i.d. Gaussian noise is added to the input $X$, excluding the CBF parameters, to simulate observation and dynamics model uncertainties. Further implementation details follow [kim_learning_2025].